# Multiobjective Aerodynamic Design Optimization of the NASA Common Research Model

### v1

@@ 3.1 Multiobjective Bayesian optimization
In contrast to single objective BO, here we first fit $K$ independent posterior GP models: $\mathcal{M}_{1},\ldots,\mathcal{M}_{K}$ for the $K$ objectives, with observations $\boldsymbol{\mathcal{D}}^{1:K}_{n}\triangleq\{\boldsymbol{\mathcal{D}}_{n}^{1},\ldots,\boldsymbol{\mathcal{D}}_{n}^{K}\}$. In the multiobjective BO, acquisition functions are typically defined in terms of a hypervolume (HV) indicator. HV of a Pareto frontier $\mathcal{Y}^{*}$ is, loosely speaking, the volume bounded by the Pareto frontier and a reference point $\mathbf{r}\in\mathbb{R}^{K}$; $\mathbf{r}$ is typically chosen as the independent infimums of each objective (in the maximization context). Expected hypervolume improvement (EHVI) [10, 19] computes the expectation (with respect to the posterior GP) of the improvement in HV due to a candidate point. The HV computation is nontrivial and scales poorly with number of objectives. Our approach, however, does not involve the computation of HV but is a direct extension of Equation 3 to the multiobjective setting.
3.2 $q$POTS: Batch Pareto optimal Thompson sampling for MOBO
Inspired by Thompson sampling (TS), our method—batch Pareto optimal Thompson sampling ($q\texttt{POTS}$[43]—selects points based on the probability that they are Pareto optimal. Specifically, we sample $\mathbf{x}$ according to the probability
[…]
@@ 3.2 $q$POTS: Batch Pareto optimal Thompson sampling for MOBO
$\mathbf{x}_{n+1}\in\arg\max_{\mathbf{x}\in\boldsymbol{\mathcal{X}}}\{Y_{1}(\mathbf{x},\omega),\ldots,Y_{K}(\mathbf{x},\omega)\}.$
@@
$X^{*}=\arg\max_{\mathbf{x}\in\boldsymbol{\mathcal{X}}}\{Y_{1}(\mathbf{x},\omega),\ldots,Y_{K}(\mathbf{x},\omega)\},$ (5)
which can be tackled with any multiobjective solver. We use evolutionary methods (e.g., NSGA-II [17]) to combine the sample efficiency of MOBO with the global search capabilities of evolutionary algorithms.
From the predicted Pareto set $X^{*}$, we then select a batch of $q$ points $\mathbf{x}_{n+1:q}\subset X^{*}$. Each point in this batch is Pareto optimal with respect to the sampled posterior paths. Although any point from $X^{*}$ is valid in principle, we promote diversity by selecting acquisitions according to a maximin distance criterion [26, 49]. This encourages a natural trade-off between exploration and exploitation.
@@
While Equation 6 may appear computationally demanding, it is efficient in practice. This is because the Pareto set $X^{*}$ returned by evolutionary solvers has finite cardinality. As a result, we only need to compute an $N^{*}\times n$ pairwise distance matrix (where $N^{*}=|X^{*}|$) and sort the resulting vector of column-wise minima. An ascending sort enables the selection of the $q$ most diverse acquisitions as the final $q$ elements of this sorted vector.
@@ 3.3 Nyström approximation for scalability in high dimensions
In high dimensions, the posterior sampling in $q\texttt{POTS}$ becomes computationally expensive. We address this via a Nyström approximation [57] with GP posteriors. We begin by writing the reparametrization technique for a GP posterior:
@@
Therefore, we have the Nyström approximation of the posterior covariance matrix square root as $\Sigma(X)^{1/2}\approx\left(\Sigma_{mN}\Sigma_{mm}^{-1/2}\right).$ The approximation in Equation 7 requires an $m\times m$ matrix square root which costs $\mathcal{O}(m^{3})$ and an additional $\mathcal{O}(Nm^{2})$ for the matrix multiplication. Therefore, an overall cost of $\mathcal{O}(m^{3}+Nm^{2})$ is much better than $\mathcal{O}(N^{3})$, provided $m\ll N$. The choice of $X_{m}$ is an open research question; in this work, we choose them to be the subset of nondominated points in $\{X_{n},\mathbf{y}_{n}\}$ which works uniformly well in practice.
In summary, the core idea of our method is inspired by Thompson sampling in single-objective optimization: while TS selects candidate points with high probability of being optimal, our approach selects points based on their probability of being Pareto optimal. $q\texttt{POTS}$  generates acquisitions by solving a computationally inexpensive multiobjective optimization problem on posterior GP sample paths. The resulting acquisitions are selected from the predicted Pareto set such that they maximize the minimum distance to previously observed points, thereby promoting diversity and balancing exploration with exploitation. A key advantage of our approach is that selecting a batch of $q$ points incurs no additional cost compared to sequential acquisitions. Simplicity and efficiency are the central strengths of our method. Our overall methodology is summarized in Algorithm 1, and our software implementation is publicly available at https://github.com/csdlpsu/qpots.
1  Input: With data sets $\boldsymbol{\mathcal{D}}_{n}^{1:K}$, fit $K$ GP models $\mathcal{M}_{1}$ through $\mathcal{M}_{K}$ and GP hyperparameters $\boldsymbol{\Omega}_{1},\ldots,\boldsymbol{\Omega}_{K}$.
@@
     3. Solve the cheap multiobjective optimization problem Equation 5 via evolutionary algorithms.
@@
     5. Observe oracles and constraints and append to data set $\boldsymbol{\mathcal{D}}^{1:K}_{i}$.
@@ 3.4 Related work
Naturally, one may seek to extend ideas from single-objective Bayesian optimization (BO) to the multiobjective setting; indeed, this intuition underlies much of the development in multiobjective Bayesian optimization (MOBO). Among the most prevalent approaches are those based on scalarization, wherein the $K$ objectives are combined into a single scalar objective, allowing the application of standard single-objective BO methods. For instance, multiobjective efficient global optimization (ParEGO) [31] performs scalarization using randomly drawn weights and applies expected improvement (EI) to the resulting objective. Multiobjective evolutionary algorithm based on decomposition (MOEA/D-EGO) [62] extends ParEGO to batch settings by employing multiple scalarizations and optimizing them in parallel via a genetic algorithm [64]. Other extensions include qParEGO and its noisy counterpart, qNParEGO [14], designed for batch and noisy evaluations, respectively. Paria et al. [36] similarly apply Chebyshev scalarization but use Thompson sampling (TS) to optimize the scalarized objective. Zhang et al. [63] propose a hypervolume-based scalarization whose expected value, under a specific distribution over weights, is equivalent to the hypervolume indicator. Their method adopts an upper confidence bound framework, although it is evaluated only on limited benchmarks. Despite their simplicity, scalarization-based approaches tend to be sample inefficient and often struggle with non-convex, disconnected, or otherwise complex Pareto frontiers.
@@
Beyond scalarization, many MOBO methods directly construct acquisition functions based on the hypervolume indicator. These approaches aim to quantify expected improvement in hypervolume, but doing so typically leads to non-convex optimization problems that are difficult to solve. The expected hypervolume improvement (EHVI) [10, 19] is a canonical extension of EI to the multiobjective setting, with several gradient-based variants also available [61, 12]. Likewise, the stepwise uncertainty reduction (SUR) criterion has been adapted for MOBO [38]. More recent methods [13, 14] extend EHVI to batch acquisitions ($q>1$) with differentiable formulations to facilitate gradient-based optimization. Nevertheless, these approaches often suffer from noisy acquisition functions and increasing optimization difficulty as batch size grows.
@@
Recognizing the challenges posed by the non-convexity of hypervolume-based acquisition functions, several recent methods introduce heuristics to better balance exploration and exploitation. Diversity-Guided Efficient Multiobjective Optimization (DGEMO) [32] selects diverse candidates by solving local optimization problems that improve hypervolume, and scales well to large batch sizes. However, DGEMO does not handle noisy observations and lacks support for constraints. Pareto Active Learning (PAL) [66] takes a classification-based approach by assigning to each candidate the probability of being Pareto optimal, based on GP posterior uncertainty. The resulting acquisition function is again difficult to optimize and is typically addressed by discretizing the design space [66].
@@
An alternative to hypervolume improvement is offered by entropy-based methods, which quantify expected reduction in uncertainty of the Pareto frontier. Predictive entropy search for multiobjective optimization (PESMO) [23] and Max-value Entropy Search for Multiobjective optimization (MESMO) [2] are two prominent examples. MESMO improves on PESMO in scalability and computational cost, but still lacks a closed-form acquisition function and depends on sampling approximations. Pareto Frontier Entropy Search (PFES) [50] offers a closed-form entropy criterion via random Fourier features (RFFs), but is restricted to stationary GP kernels. Similarly, Joint Entropy Search for Multiobjective Optimization (JESMO) [25, 52] estimates joint information gain over inputs and outputs, but still requires approximations to the acquisition function.
@@
Despite these developments, the use of TS in MOBO remains relatively unexplored. Existing methods that incorporate TS often do so in conjunction with traditional scalarization or hypervolume techniques, inheriting their associated limitations. For example, Thompson Sampling Efficient Multiobjective Optimization (TSEMO) [6] samples Pareto sets from GP posteriors and evaluates them using hypervolume improvement. TS with Chebyshev scalarization (TS-TCH) likewise depends on scalarized objectives. Multiobjective Regionalized Bayesian Optimization (MORBO) [15] introduces a trust-region strategy and uses TS to identify local candidates; however, it ultimately relies on hypervolume improvement for selection and its capability to handle constraints remains unclear.
@@ 4 Experiments
We now demonstrate our proposed approach, $q\texttt{POTS}$, on several synthetic and real-world experiments of varying input ($d$) and output ($K$) dimensions. We compare $q\texttt{POTS}$ against TSEMO (another popular TS-based approach), PESMO, JESMO, MESMO (entropy-based approaches), qNEHVI, qPAREGO (classical approaches), and random (Sobol) sampling.
@@ 4.1 Synthetic experiments
We consider the following synthetic experiments: the Branin-Currin function, the ZDT3 function which has disconnected Pareto frontiers, the DTLZ3 function, and the DTLZ7 test function. We provide each experiment with a set of $n=10\times d$ seed samples and observations to start the algorithm, chosen uniformly at random from $\mathcal{X}$; then they are repeated $10$ times. However, the seed is fixed for each repetition across all acquisition policies to keep the comparison fair. We add Gaussian noise with variance $\tau^{2}=10^{-3}$ to all our observations to assist with the conditioning of the GP covariance matrix. The experiments are run with a high-performance compute node with 256GB memory and $48$-core nodes; repetitions are parallelized across the cores. Our metric for comparison is the hypervolume indicator computed via box decomposition [60].
Throughout the manuscript, we default to the anisotropic Matérn class kernel with $\nu=5/2$. Our implementation primarily builds on GPyTorch [22] and BoTorch [1]; we leverage MPI4Py [11] to parallelize our repetitions. We use PyMOO [3] to leverage its NSGA-II solver in our inner optimization (Equation 5); the population size for all problems is fixed at $100\times d$. The full source code is available at https://github.com/csdlpsu/qpots.
Figure 3 shows the performance comparison (in terms of hypervolume) for sequential sampling ($q=1$). Notice that $q\texttt{POTS}$ (blue) is consistently the best performer, or is amongst the best. We show $q\texttt{POTS}$-Nystrom-Pareto (that is $X_{m}$ chosen as the current Pareto set) for the $d=10$ experiments which shows no loss of accuracy for the DTLZ3 function and, surprisingly, a slight gain in accuracy for the ZDT3 function. Despite the success in the $q=1$ setting, our main benefit is seen in the batch sampling setting which we discuss next.
@@
Figure 3: Hypervolume (higher the better) Vs. iterations for sequential ($q=1$) acquisition; plots show mean and ±$\pm 1$ standard deviation out of $10$ repetitions.
The demonstration for the batch setting is shown in Figure 4, where we present experiments with $q=2,4$; like in the $q=1$ setting we include $q\texttt{POTS}$-Nystrom-Pareto for the $d=10$ experiments only. While $q\texttt{POTS}$ comfortably wins in all experiments, notice that the difference with TSEMO is more pronounced in the batch setting. The advantage in the batch setting $q>1$ is a key benefit of $q\texttt{POTS}$, because the cost of batch sampling is practically the same as sequential ($q=1$) sampling.
@@
Figure 4: Batch acquisition. Hypervolume Vs. iterations for batch ($q>1$) acquisition; plots show mean and ±$\pm 1$ standard deviation out of $10$ repetitions. $q\texttt{POTS}$ outperforms all competitors, but the benefit is more pronounced in the batch case. Additional experiments, including constrained problems, are shown in the supplementary material.

### v2

@@ 3.1 Multiobjective Bayesian optimization
In contrast to single objective BO, here we first fit $K$ independent posterior GP models: $\mathcal{M}_{1},\ldots,\mathcal{M}_{K}$ for the $K$ objectives, with observations $\boldsymbol{\mathcal{D}}^{1:K}_{n}\triangleq\{\boldsymbol{\mathcal{D}}_{n}^{1},\ldots,\boldsymbol{\mathcal{D}}_{n}^{K}\}$. In the multiobjective BO, acquisition functions are typically defined in terms of a hypervolume (HV) indicator. HV of a Pareto frontier $\mathcal{Y}^{*}$ is, loosely speaking, the volume bounded by the Pareto frontier and a reference point $\mathbf{r}\in\mathbb{R}^{K}$; $\mathbf{r}$ is typically chosen as the independent infimums of each objective (in the maximization context). Expected hypervolume improvement (EHVI) [couckuytFastCalculationMultiobjective2014, emmerichComputationExpectedImprovement2008] computes the expectation (with respect to the posterior GP) of the improvement in HV due to a candidate point. The HV computation is nontrivial and scales poorly with number of objectives. Our approach, however, does not involve the computation of HV but is a direct extension of Equation˜3 to the multiobjective setting.
3.2 qqPOTS: Batch Pareto optimal Thompson sampling for MOBO
Inspired by Thompson sampling (TS), our method—batch Pareto optimal Thompson sampling ($q\texttt{POTS}$[renganathan2025q]—selects points based on the probability that they are Pareto optimal. Specifically, we sample $\mathbf{x}$ according to the probability
$\begin{split}p_{\mathcal{X}^{*}}(\mathbf{x})=\int\delta\left(\mathbf{x}-\arg\max_{\mathbf{x}\in\mathcal{X}}\{Y_{1}(\mathbf{x}),\ldots,Y_{K}(\mathbf{x})\}\right)\,p(Y_{1}\mid\boldsymbol{\mathcal{D}}^{1}_{n})\cdots p(Y_{K}\mid\boldsymbol{\mathcal{D}}^{K}_{n})\,dY_{1}\cdots dY_{K},\end{split}$ (4)
@@
$\mathbf{x}_{n+1}\in\arg\max_{\mathbf{x}\in\mathcal{X}}\{Y_{1}(\mathbf{x},\omega),\ldots,Y_{K}(\mathbf{x},\omega)\}.$
@@
$X^{*}=\arg\max_{\mathbf{x}\in\mathcal{X}}\{Y_{1}(\mathbf{x},\omega),\ldots,Y_{K}(\mathbf{x},\omega)\},$ (5)
which can be tackled with any multiobjective solver. We use evolutionary methods (e.g., NSGA-II [deb2002fast]) to combine the sample efficiency of MOBO with the global search capabilities of evolutionary algorithms.
From the predicted Pareto set $X^{*}$, we then select a batch of $q$ points $\mathbf{x}_{n+1:q}\subset X^{*}$. Each point in this batch is Pareto optimal with respect to the sampled posterior paths. Although any point from $X^{*}$ is valid in principle, we promote diversity by selecting acquisitions according to a maximin distance criterion [johnson1990minimax, sun2019synthesizing]. This encourages a natural trade-off between exploration and exploitation.
@@
While Equation˜6 may appear computationally demanding, it is efficient in practice. This is because the Pareto set $X^{*}$ returned by evolutionary solvers has finite cardinality. As a result, we only need to compute an $N^{*}\times n$ pairwise distance matrix (where $N^{*}=|X^{*}|$) and sort the resulting vector of column-wise minima. An ascending sort enables the selection of the $q$ most diverse acquisitions as the final $q$ elements of this sorted vector.
@@ 3.3 Nyström approximation for scalability in high dimensions
In high dimensions, the posterior sampling in $q\texttt{POTS}$ becomes computationally expensive. We address this via a Nyström approximation [williamsUsingNystromMethod2000] with GP posteriors. We begin by writing the reparametrization technique for a GP posterior:
@@
Therefore, we have the Nyström approximation of the posterior covariance matrix square root as $\Sigma(X)^{1/2}\approx\left(\Sigma_{mN}\Sigma_{mm}^{-1/2}\right).$ The approximation in Equation˜7 requires an $m\times m$ matrix square root which costs $\mathcal{O}(m^{3})$ and an additional $\mathcal{O}(Nm^{2})$ for the matrix multiplication. Therefore, an overall cost of $\mathcal{O}(m^{3}+Nm^{2})$ is much better than $\mathcal{O}(N^{3})$, provided $m\ll N$. The choice of $X_{m}$ is an open research question; in this work, we choose them to be the subset of nondominated points in $\{X_{n},\mathbf{y}_{n}\}$ which works uniformly well in practice.
In summary, the core idea of our method is inspired by Thompson sampling in single-objective optimization: while TS selects candidate points with high probability of being optimal, our approach selects points based on their probability of being Pareto optimal. $q\texttt{POTS}$  generates acquisitions by solving a computationally inexpensive multiobjective optimization problem on posterior GP sample paths. The resulting acquisitions are selected from the predicted Pareto set such that they maximize the minimum distance to previously observed points, thereby promoting diversity and balancing exploration with exploitation. A key advantage of our approach is that selecting a batch of $q$ points incurs no additional cost compared to sequential acquisitions. Simplicity and efficiency are the central strengths of our method. Our overall methodology is summarized in Algorithm˜1, and our software implementation is publicly available at https://github.com/csdlpsu/qpots.
1Input: With data sets $\boldsymbol{\mathcal{D}}_{n}^{1:K}$, fit $K$ GP models $\mathcal{M}_{1}$ through $\mathcal{M}_{K}$ and GP hyperparameters $\boldsymbol{\Omega}_{1},\ldots,\boldsymbol{\Omega}_{K}$.
@@
  3. Solve the cheap multiobjective optimization problem Equation˜5 via evolutionary algorithms.
@@
  5. Evaluate objectives and append to data set $\boldsymbol{\mathcal{D}}^{1:K}_{i}$.
@@ 3.4 Related work
Naturally, one may seek to extend ideas from single-objective Bayesian optimization (BO) to the multiobjective setting; indeed, this intuition underlies much of the development in multiobjective Bayesian optimization (MOBO). Among the most prevalent approaches are those based on scalarization, wherein the $K$ objectives are combined into a single scalar objective, allowing the application of standard single-objective BO methods. For instance, multiobjective efficient global optimization (ParEGO) [knowles2006parego] performs scalarization using randomly drawn weights and applies expected improvement (EI) to the resulting objective. Multiobjective evolutionary algorithm based on decomposition (MOEA/D-EGO) [zhang2009expensive] extends ParEGO to batch settings by employing multiple scalarizations and optimizing them in parallel via a genetic algorithm [zhou2012multiobjective]. Other extensions include qParEGO and its noisy counterpart, qNParEGO [daultonParallelBayesianOptimization2021], designed for batch and noisy evaluations, respectively. Paria et al. [paria2020flexible] similarly apply Chebyshev scalarization but use Thompson sampling (TS) to optimize the scalarized objective. Zhang et al. [zhang2020random] propose a hypervolume-based scalarization whose expected value, under a specific distribution over weights, is equivalent to the hypervolume indicator. Their method adopts an upper confidence bound framework, although it is evaluated only on limited benchmarks. Despite their simplicity, scalarization-based approaches tend to be sample inefficient and often struggle with non-convex, disconnected, or otherwise complex Pareto frontiers.
@@
Beyond scalarization, many MOBO methods directly construct acquisition functions based on the hypervolume indicator. These approaches aim to quantify expected improvement in hypervolume, but doing so typically leads to non-convex optimization problems that are difficult to solve. The expected hypervolume improvement (EHVI) [couckuytFastCalculationMultiobjective2014, emmerichComputationExpectedImprovement2008] is a canonical extension of EI to the multiobjective setting, with several gradient-based variants also available [yang2019multi, daulton2020differentiable]. Likewise, the stepwise uncertainty reduction (SUR) criterion has been adapted for MOBO [picheny2015multiobjective]. More recent methods [daultonDifferentiableExpectedHypervolume2020, daultonParallelBayesianOptimization2021] extend EHVI to batch acquisitions ($q>1$) with differentiable formulations to facilitate gradient-based optimization. Nevertheless, these approaches often suffer from noisy acquisition functions and increasing optimization difficulty as batch size grows.
@@
Recognizing the challenges posed by the non-convexity of hypervolume-based acquisition functions, several recent methods introduce heuristics to better balance exploration and exploitation. Diversity-Guided Efficient Multiobjective Optimization (DGEMO) [konakoviclukovicDiversityGuidedMultiObjectiveBayesian2020] selects diverse candidates by solving local optimization problems that improve hypervolume, and scales well to large batch sizes. However, DGEMO does not handle noisy observations. Pareto Active Learning (PAL) [zuluaga2013active] takes a classification-based approach by assigning to each candidate the probability of being Pareto optimal, based on GP posterior uncertainty. The resulting acquisition function is again difficult to optimize and is typically addressed by discretizing the design space [zuluaga2013active].
@@
An alternative to hypervolume improvement is offered by entropy-based methods, which quantifies expected reduction in uncertainty of the Pareto frontier. Predictive entropy search for multiobjective optimization (PESMO) [hernandez-lobatoPredictiveEntropySearch] and Max-value Entropy Search for Multiobjective optimization (MESMO) [belakariaMaxvalueEntropySearch2019] are two prominent examples. MESMO improves on PESMO in scalability and computational cost, but still lacks a closed-form acquisition function and depends on sampling approximations. Pareto Frontier Entropy Search (PFES) [suzuki2020multi] offers a closed-form entropy criterion via random Fourier features (RFFs), but is restricted to stationary GP kernels. Similarly, Joint Entropy Search for Multiobjective Optimization (JESMO) [hvarfner2022joint, tu2022joint] estimates joint information gain over inputs and outputs, but still requires approximations to the acquisition function.
@@
Despite these developments, the use of TS in MOBO remains relatively unexplored. Existing methods that incorporate TS often do so in conjunction with traditional scalarization or hypervolume techniques, inheriting their associated limitations. For example, Thompson Sampling Efficient Multiobjective Optimization (TSEMO) [bradford2018efficient] samples Pareto sets from GP posteriors and evaluates them using hypervolume improvement. TS with Chebyshev scalarization (TS-TCH) likewise depends on scalarized objectives. Multiobjective Regionalized Bayesian Optimization (MORBO) [daulton2022multi] introduces a trust-region strategy and uses TS to identify local candidates; however, it ultimately relies on hypervolume improvement for selection.
@@
The advantages of treating multipoint design as a multi-objective problem should be more explicitly articulated: why and when Pareto sets are preferable to weighted-sum scalarization in aerospace contexts. - Provide a stronger aerodynamic interpretation of the Pareto front in the CRM case to demonstrate the practical value of multiple solutions, not just numerical improvement.
@@ 4 Experiments
We now demonstrate our proposed approach, $q\texttt{POTS}$, on several synthetic and real-world experiments of varying input ($d$) and keeping output ($K$) dimensions to $K=2$, to keep it consistent with the CRM experiment although our method seamlessly extends to any $K$. We compare $q\texttt{POTS}$ against TSEMO (another popular TS-based approach), PESMO, JESMO, MESMO (entropy-based approaches), qNEHVI, qPAREGO (classical approaches), and random (Sobol) sampling.
Experiment Branin-Currin DTLZ3 DTLZ7 ZDT3 NASA CRM
Reference point $(70,70)$ $(10^{4},10^{4})$ $(25,25)$ $(11,11)$ $(10^{2},10^{2})$
Dimensionality $(d,K)$ $(2,2)$ $(10,2)$ $(5,2)$ $(10,2)$ $(24,2)$
Seed/repetitions $20/10$ $100/10$ $50/10$ $100/10$ $200/1$
Batch size ($q$) $1,4$ $1,4$ $1,4$ $1,4$ 11
Table 1: Summary of experiment settings.
@@ 4.1 Synthetic experiments
We consider the following synthetic experiments: the Branin-Currin function, the ZDT3 function, which has disconnected Pareto frontiers, the DTLZ3 function, and the DTLZ7 test function. We provide each experiment with a set of $n=10\times d$ seed samples and observations to start the algorithm, chosen uniformly at random from $\mathcal{X}$; then they are repeated $10$ times. However, the seed is fixed for each repetition across all acquisition policies to keep the comparison fair. We add Gaussian noise with variance $\tau^{2}=10^{-3}$ to all our observations to assist with the conditioning of the GP covariance matrix. The experiments are run with a high-performance compute node with 256GB memory and $48$-core nodes; repetitions are parallelized across the cores. Our metric for comparison is the hypervolume indicator computed via box decomposition [yang2019efficient]. It should be noted that hypervolume computation is a postprocessing step in the methodology and hence is not part of the algorithm. The hypervolume computation involves specifying a reference point (which is the absolute worst value each objective can take) – these are summarized in Table˜1. For synthetic experiment, we determine this by densely sampling the input domain $\mathcal{X}$ and evaluating the objectives on these samples. For the NASA CRM experiment, we use domain knowledge (that drag coefficients are lower bounded by 0 and hence set $(100,100)$.
Throughout the manuscript, we default to the anisotropic Matérn class kernel with $\nu=5/2$. Our implementation primarily builds on GPyTorch [gardner2018gpytorch] and BoTorch [balandat2019botorch]; we leverage MPI4Py [dalcin2021mpi4py] to parallelize our repetitions. We use PyMOO [blank2020pymoo] to leverage its NSGA-II solver in our inner optimization (Equation˜5); the population size for all problems is fixed at $100\times d$.
Figure˜4 shows the performance comparison (in terms of hypervolume) for sequential sampling ($q=1$). Notice that $q\texttt{POTS}$ (blue) is consistently the best performer, or is amongst the best. We show $q\texttt{POTS}$-Nystrom-Pareto (that is $X_{m}$ chosen as the current Pareto set) for the $d=10$ experiments which shows no loss of accuracy for the DTLZ3 function and, surprisingly, a slight gain in accuracy for the ZDT3 function. Despite the success in the $q=1$ setting, our main benefit is seen in the batch sampling setting which we discuss next.
@@
Figure 4: Hypervolume (higher the better) Vs. iterations for sequential ($q=1$) acquisition; plots show mean and $\pm 1$ standard deviation out of $10$ repetitions.
The demonstration for the batch setting is shown in Figure˜5, where we present experiments with $q=2,4$; like in the $q=1$ setting we include $q\texttt{POTS}$-Nystrom-Pareto for the $d=10$ experiments only. While $q\texttt{POTS}$ comfortably wins in all experiments, notice that the difference with TSEMO is more pronounced in the batch setting. The advantage in the batch setting $q>1$ is a key benefit of $q\texttt{POTS}$, because the cost of batch sampling is practically the same as sequential ($q=1$) sampling. It is expected that $q\texttt{POTS}$ performs better in the batch setting; this is for the following reasons. While the Thompson sampling automatically balances exploration and exploitation in the iterates, we further ensure diversity in the candidates by sorting them according to the maximin distance. Additionally, our batch sampling avoids complex stochastic optimization as in other methods (e.g., qNEHVI) which makes the computation cheap and deterministic.
@@
Figure 5: Batch acquisition. Hypervolume Vs. iterations for batch ($q>1$) acquisition; plots show mean and $\pm 1$ standard deviation out of $10$ repetitions. $q\texttt{POTS}$ outperforms all competitors, but the benefit is more pronounced in the batch case.